### v1

@@ Abstract.
We find the shortest realized stretch factor for a fully irreducible $\varphi\in\textup{Out}(F_{3})$ and show that it is realized by a “principal” fully irreducible element. We also show that it is the only principal fully irreducible produced by a single fold in any rank.
@@ 1. Introduction
Let $F_{r}$ denote the free group of rank $r\geq 2$, and consider its outer automorphism group $\textup{Out}(F_{r})$. Each element of $\textup{Out}(F_{r})$ has an associated stretch factor (also called growth rate or dilatation)
@@
Here $\|w\|$ denotes the cyclically reduced word length with respect to some fixed basis. The stretch factor records how fast elements grow under iteration of $\varphi$, and is independent of the chosen basis and the representative of the outer class.
For a fully irreducible element, where no power $\varphi^{k}$ preserves a proper free factor, Bestvina and Handel [BH92] construct an “irreducible train track representative” for $\varphi$, a self homotopy equivalence of a graph with good behavior under iteration which induces $\varphi$ on the fundamental group. They show that the stretch factor of $\varphi$ is realised as the Perron–Frobenius eigenvalue of the transition matrix, a non-negative integer matrix associated to the train track map.
@@
The ingredients in the Theorem A proof are Lemma 5.1, showing stretch factors of fully irreducible elements are Perron numbers, and verification that $\mathfrak{g}$ is an irreducible train track representative of a fully irreducible outer automorphism realising the minimum achievable Perron number.
@@
We also prove the map $\psi$ is principal in the sense of [AKKP19], where it is proved they possess a stability property mimicking that held by pseudo-Anosov mapping classes and used in [KMPT22b] and [KMPT22a] to understand a typical outer automorphism and tree in $\partial\textup{CV}_{r}$. Their action on Culler–Vogtmann Outer space generally more closely resembles a hyperbolic setting: While all fully irreducible elements act loxodromically, they only have an axis bundle rather than a unique axis [HM11]. But principal fully irreducible outer automorphisms do have a unique axis, reflecting the uniqueness of the Stallings fold decompositions of their irreducible train track representatives. This axis passes through the highest dimensional simplices of Outer space.
@@
The name is chosen to reflect similarities with the principal pseudo-Anosov mapping classes of a surface, as used for instance by Masur in [Mas82]. Every pseudo-Anosov mapping class has a representative leaving invariant a pair of transverse measured singular minimal foliations, one of which is expanded and the other contracted by the homeomorphism. Pseudo-Anosovs act loxodromically on Teichmüller space, and these two foliations provide the endpoints of its axis in the Thurston boundary. A pseudo-Anosov is said to be principal if these foliations have only 3-pronged singularities. Gadre and Maher [GM17] proved principal pseudo-Anosov mapping classes random-walk generic. The “singularity structure” of a fully irreducible outer automorphism, namely its ideal Whitehead graph, is similarly controlled by its attracting endpoint, and in the case of a principal element will be a union of the maximal number of triangles. As such, principal elements are the subset of those proved generic in [KMPT22b] that possess the stability property of [AKKP19].
@@
Given a train track map, we can factor it as a series of folds followed by a homeomorphism. Up to a reasonable equivalence relation, the number of folds in this decomposition is an invariant of a principal fully irreducible outer automorphism, and so we can try to minimise it. It turns out that the same automorphism achieves the minimum complexity — over all ranks — in this sense.
@@
The tool we use for characterising rank-3 principal fully irreducible outer automorphisms is a directed graph $\widehat{\mathcal{A}_{3}}$ we call the principal stratum automaton3). This is a refinement of the lonely direction automaton of [GP23], which was used to determine the graphs carrying train track representatives of principal fully irreducible elements of $\textup{Out}(F_{3})$. That theorem was proved by passing to a rotationless power of $\varphi$, which is unhelpful for minimising other kinds of complexity such as the stretch factor or length of a Stallings fold decomposition. Our refinement allows us to work with the automorphism as it comes, allowing for the periodic behavior absent in the rotationless case.
@@
In the process, we also prove in Proposition 3.4 that any principal axis in $\textup{CV}_{3}$ passes through the $\textup{Out}(F_{3})$ orbit of simplices given by the graph shown above.
@@ Acknowledgements
We are grateful to the Women in Groups Geometry and Dynamics (WiGGD) program from which this paper arose, and to Anna Parlak for discussions on related questions. We are grateful to Lee Mosher for providing his talents as a rabbit-hole preventing sounding board, and to Ilya Kapovich for helping us track down a reference and his ongoing interest in our work. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 850930), an NSF postdoctoral fellowship, and an NSERC Discovery Grant.
@@ 2. Background
Assume throughout this section $\Gamma$ is a finite oriented graph where each vertex has valence at least $3$ and $F_{r}$ is a free group of rank $r\geq 3$.
@@ 2.1. Edge Maps on Graphs
Suppose $\Gamma$ has positively oriented edges $\{e_{1},e_{2}\dots,e_{n}\}$ and vertices $\{v_{1},v_{2},\dots,v_{m}\}$. We use the notation $E\Gamma:=\{e_{1},e_{2}\dots,e_{n}\}$, and $V\Gamma:=\{v_{1},v_{2},\dots,v_{m}\}$, and $E^{\pm}\Gamma=\{e_{1},\overline{e_{1}},\dots,e_{n},\overline{e_{n}}\}$, with an overline indicating a reversal of orientation.
Given $v\in V\Gamma$, a direction at $v$ will mean an element of $E^{\pm}\Gamma$ with initial vertex $v$. We let $\mathcal{D}\Gamma$ denote the set of directions at vertices in $\Gamma$. A turn at $v$ will mean an unordered pair $\{d_{1},d_{2}\}$ of directions at $v$. The turn is degenerate if $d_{1}=d_{2}$.
[Uncaptioned image]An edge path (or just path) $\rho$ in $\Gamma$ is a finite sequence $(a_{1},a_{2},\dots,a_{\ell})\in(E^{\pm}\Gamma)^{\ell}$ such that there exists a sequence $(v_{1},v_{2},\dots,v_{\ell-1})\in(V\Gamma)^{\ell-1}$ satisfying that the turn $\{\overline{a_{j}},a_{j+1}\}$ is a turn at $v_{j}$ for each $j\in\{1,2,\dots,\ell-1\}$. For such a path $(a_{1},a_{2},\dots,a_{m})$ we write $\gamma=a_{1}a_{2}\dots a_{m}$ and say $\gamma$ contains the oriented edges $a_{1},a_{2},\dots,a_{m}$ and takes the turns $\{\overline{a_{1}},a_{2}\}$, $\{\overline{a_{2}},a_{3}\}$, $\dots$, $\{\overline{a_{n-1}},a_{n}\}$. We call $\gamma$ tight if it takes no degenerate turns, which we colloquially describe as there being no “backtracking.”
@@
To $g$ we associate a direction map $Dg:\mathcal{D}\Gamma\to\mathcal{D}\Gamma^{\prime}$ such that if $g(e)=a_{1}a_{2}\dots a_{m}$, for some $m\geq 1$ and $a_{1},a_{2},\dots,a_{m}\in E^{\pm}\Gamma$, then $Dg(e)=a_{1}$. We call a direction $e$ periodic if $Dg^{k}(e)=e$ for some $k>0$, and fixed if $k=1$. When $g\colon\Gamma\to\Gamma$ is a self-map, the turn $\{d_{1},d_{2}\}$ is called an illegal turn for $g$ if $\{Dg^{k}(d_{1}),Dg^{k}(d_{2})\}$ is degenerate for some $k$. Defining an equivalence relation on $\mathcal{D}\Gamma$ by $d_{1}\sim d_{2}$ when $\{d_{1},d_{2}\}$ is an illegal turn, the equivalence classes are called gates. Note that each gate contains a unique periodic direction.
Viewing $g$ as a continuous map of graphs, we say $g$ represents $\varphi$ when $\pi_{1}(\Gamma)$ has been identified with $F_{r}$ (that is, $\Gamma$ is marked) and $\varphi$ is the induced map of fundamental groups. When a marking is not explicitly given, we mean that “there exists a marking such that.”
@@ 2.2. Train track maps and fully irreducible outer automorphisms
A nonnegative integral matrix $A=[a_{ij}]$ is irreducible if for each $(i,j)$, there is a $k\in\mathbb{Z}_{>0}$ so that the ${ij}^{th}$ entry of $A^{k}$ is positive, and so in particular at least $1$. If $A^{k}$ is strictly positive for some $k\in\mathbb{Z}_{>0}$ then $A$ is primitive. Further, $A$ is Perron–Frobenius (PF) if there exists an $N$ such that, for each $k\geq N$, we have that $A^{k}$ is strictly positive.
For nonnegative integral matrices, being primitive is equivalent to being irreducible aperiodic. If $M=M(g)$ is primitive, then by Perron–Frobenius theory, $M$ has a unique eigenvalue $\lambda(g)$ of maximal modulus, $\lambda(g)$ is real, and $\lambda(g)>1$. This $\lambda(g)$ is the Perron–Frobenius (PF) eigenvalue of $M$ and is called the stretch factor (or dilatation) of $g$.
A tt map is irreducible if its transition matrix is irreducible. Not every element of $\textup{Out}(F_{r})$ is represented by a train track map, and even fewer by irreducible train track maps. An outer automorphism $\varphi\in\textup{Out}(F_{r})$ is fully irreducible if no positive power preserves the conjugacy class of a proper free factor of $F_{r}$. Bestvina and Handel [BH92] proved that every fully irreducible outer automorphism admits expanding irreducible train track representatives and that the stretch factor of these representatives is an invariant of the outer automorphism.
By [FH11, Corollary 4.43], for each $r\geq 2$, there exists an $R(r)\in\mathbb{N}$ such that for each expanding irreducible train track representative $g$ of a fully irreducible $\varphi\in\textup{Out}(F_{r})$, among other properties, each periodic direction is fixed by $g^{R(r)}$. This power $R$ is called the rotationless power.
@@ 2.3. Whitehead graphs & lamination train track (ltt) structures
Let $g:\Gamma\to\Gamma$ be a tt map. The local Whitehead graph $\operatorname{LW}(g;v)$ at a $v\in V\Gamma$ has a vertex for each direction at $v$ and edge connecting the vertices corresponding to a pair of directions $\{d_{1},d_{2}\}$ at $v$ precisely when the turn $\{d_{1},d_{2}\}$ is $g^{k}$-taken for some $k\in\mathbb{Z}_{>0}$. Given a fixed vertex $v$, the stable Whitehead graph $\operatorname{SW}(g;v)$ is the restriction of $\operatorname{LW}(g;v)$ to the periodic direction vertices and the edges between them. In terms of gates, $\operatorname{SW}(g;v)$ has a vertex for each gate at $v$.
@@
$\operatorname{IW}(\varphi)\cong\bigsqcup_{v\in V\Gamma}\operatorname{SW}(g;v),$
@@
The lamination train track (ltt) structure $G(g)$ is obtained from its underlying graph $\Gamma$ by replacing each vertex $v\in V\Gamma$ with $\operatorname{LW}(g;v)$: replace $v$ with a vertex for each directed edge at $v$ labeled with that direction. Then identify each of these new vertices with the corresponding vertex of $\operatorname{LW}(g;v)$. Vertices and edges of $\operatorname{SW}(g;v)$ are colored purple and the remaining vertices and (open) edges are colored red. Alternatively, one could start with $\bigsqcup_{v\in V\Gamma}\operatorname{LW}(g;v)$, color the $\operatorname{LW}(g;v)$ as just described, and then include a directed edge $[e,\overline{e}]$ for each directed edge $e\in E\Gamma$. To simplify figures, if the local Whitehead graph at a vertex of $\Gamma$ is complete we do not always draw it.
Figure 1. The right-hand image is the ltt structure $G(\mathfrak{g})$ for the map $\mathfrak{g}$ of (11) with the taken turns as in (12). The 3 local Whitehead graphs are colored, with the stable Whitehead graphs in purple (missing only $\bar{c}$). The ideal Whitehead graph is the union of the purple graphs, i.e. of the stable Whitehead graphs.

### v2

@@ Abstract.
We find the shortest realized stretch factor for a fully irreducible $\varphi\in\textup{Out}(F_{3})$ and show that it is realized by a “principal” fully irreducible element. We also show that it is the only principal fully irreducible outer automorphism in any rank produced by a single fold.
@@ 1. Introduction
Let $F_{r}$ denote the free group of rank $r\geq 2$, and consider its outer automorphism group $\textup{Out}(F_{r})$. Each element of $\textup{Out}(F_{r})$ has an associated stretch factor (also called its growth rate or dilatation)
@@
Here $\|w\|$ denotes the cyclically reduced word length with respect to some fixed free basis of $F_{r}$. The stretch factor records how fast elements grow under iteration of $\varphi$, and is independent of the chosen basis and the representative of the outer class.
For a fully irreducible element, where no power $\varphi^{k}$ preserves a proper free factor, Bestvina and Handel [BH92] construct an “irreducible train track representative” for $\varphi$, a self homotopy equivalence of a graph with good behavior under iteration and which induces $\varphi$ on the fundamental group. They show that the stretch factor of $\varphi$ is realised as the Perron–Frobenius eigenvalue of the transition matrix, a non-negative integer matrix associated to the train track representative.
@@
The ingredients in the Theorem A proof are Lemma 5.1, showing stretch factors of fully irreducible outer automorphisms are Perron numbers, and verification that $\mathfrak{g}$ is an irreducible train track representative of a fully irreducible outer automorphism realising the minimum achievable Perron number.
We use a result of Boyd listing the smallest Perron numbers for several fixed algebraic degrees [Boy85]. Boyd lists the smallest (and in some degrees the second, third, or fourth smallest) Perron numbers for fixed algebraic degrees between 2 and 12. For ranks $r>3$, this strategy of checking the smallest Perron numbers with algebraic degree in a fixed range will likely be less fruitful. The known examples of fully irreducible outer automorphisms with small stretch factors are quite a bit larger than the smallest Perron numbers listed in [Boy85].
Given a train track map, we can factor it as a series of folds followed by a homeomorphism. It is notable that the train track map $\mathfrak{g}$ representing $\psi$ in Theorem A has a fold decomposition consisting of a single fold and a homeomorphism. As folds increase the number of edges in the image of a train track map, it seems reasonable to conjecture that outer automorphisms with minimal stretch factors have train track representatives with few folds in their fold decomposition. This idea is followed up and quantified in [Hil24].
In the mapping class group setting, minimal stretch factors among pseudo-Anosov elements on a fixed surface of finite type have been widely studied. Until very recently, the minimum was only known for a few isolated examples of surfaces. Consider an orientable surface of finite type with genus $g$ and $p$ punctures. Previously the minimum stretch factor among all pseudo-Anosov mapping classes was precisely known only for small values of $(g,p)$, for example $(1,0)$ and $(1,1)$, $(2,0)$ [CH08], and $(0,p)$ for $p=3,4,5,6,7,8$ ([SKL02], [HS07], [LT11b]). Recent breakthrough work of [TZ24] also computes the minimum stretch factor for $(0,p)$ for large $p$. Among orientable pseudo-Anosov mapping classes, the minimum is known for surfaces with 00 punctures and genus $g=2,3,4,5,8$ ([LT11a], [Hir10]). For surfaces with $p>0$ punctures, pseudo-Anosov mapping classes induce outer automorphisms of $F_{r}$ when $r=2g+p-1$. However, there are many outer automorphisms which do not come from surface homeomorphisms, so minimal stretch factor results in the mapping class group setting do not directly carry over to $\textup{Out}(F_{r})$.
@@
We also prove that the map $\psi$ is principal in the sense of [AKKP19], where it is also proved that they possess a stability property mimicking that held by pseudo-Anosov mapping classes and used in [KMPT22b] and [KMPT22a] to understand a typical outer automorphism and tree in $\partial\textup{CV}_{r}$. Their action on Culler–Vogtmann Outer space generally more closely resembles a hyperbolic setting: While all fully irreducible elements act loxodromically, they only have an axis bundle rather than a unique axis [HM11]. But principal fully irreducible outer automorphisms do have a unique axis, reflecting the uniqueness of the Stallings fold decompositions of their irreducible train track representatives. This axis passes through the highest dimensional simplices of Outer space.
@@
The name “principal” is chosen to reflect similarities with the principal pseudo-Anosov mapping classes of a surface, as used for example by Masur in [Mas82]. Every pseudo-Anosov mapping class has a representative leaving invariant a pair of transverse measured singular minimal foliations, one of which is expanded and the other of which is contracted by the homeomorphism. Pseudo-Anosov mapping classes act loxodromically on Teichmüller space, and the pair of transverse foliations provides the endpoints of its axis in the Thurston boundary. A pseudo-Anosov is principal if these foliations have only 3-pronged singularities. Gadre and Maher [GM17] proved principal pseudo-Anosov mapping classes random-walk generic. The “singularity structure” of a fully irreducible outer automorphism, namely its ideal Whitehead graph, is similarly controlled by its attracting endpoint, and in the case of a principal element it will be a union of the maximal number of triangles. As such, principal elements comprise the subset of those proved generic in [KMPT22b] that possess the stability property of [AKKP19].
@@
As mentioned above, each train track map can be factored as a series of Stallings folds followed by a homeomorphism. Up to a reasonable equivalence relation, the number of folds in this decomposition is an invariant of a principal fully irreducible outer automorphism, and so we can try to minimise it. It turns out that the same outer automorphism achieves the minimum complexity — over all ranks — in this sense.
@@
The tool we use for characterising rank-3 principal fully irreducible outer automorphisms is a directed graph $\widehat{\mathcal{A}_{3}}$, we call the principal stratum automaton (defined in §3). This is a refinement of the lonely direction automaton of [GP23], which was used to determine the graphs carrying train track representatives of principal fully irreducible $\varphi\in\textup{Out}(F_{3})$. That theorem was proved by passing to a rotationless power of $\varphi$, which is unhelpful for minimising other kinds of complexity such as the stretch factor or length of a Stallings fold decomposition. Our refinement allows us to work with the outer automorphism as it comes, allowing for the periodic behavior absent in the rotationless case.
@@
We note that, subsequent to the posting of this paper, Theorem C was significantly generalized in [Pfa24].
We further note that, using [MP16, Theorem 4.7], any principal fully irreducible $\varphi\in\textup{Out}(F_{r})$ has a single unique invariant axis and the single unique invariant axis of any $\psi^{-1}\circ\varphi\circ\psi$ differs from that of $\varphi\in\textup{Out}(F_{r})$ only by a change of marking. With a little work, this implies that the number of folds in a Stallings fold decompositions of the conjugacy class of a principal fully irreducible $\varphi$ is an invariant of the conjugacy class. This similarly holds for any “lone axis” fully irreducible outer automorphism. More generally, one would need to take the minimal number of folds in a fold decomposition.
In the process of proving Theorem C, we also prove in Proposition 3.4 that any principal axis in $\textup{CV}_{3}$ passes through the $\textup{Out}(F_{3})$ orbit of simplices given by the graph shown above.
@@ Acknowledgements
We are grateful to the Women in Groups Geometry and Dynamics (WiGGD) program from which this paper arose, and to Anna Parlak for discussions on related questions. We are grateful to Lee Mosher for providing his talents as a rabbit-hole preventing sounding board, and to Ilya Kapovich and Caglar Uyanik for helping us track down references and their ongoing interest in our work. We would also like to thank the referee for a careful reading. This work has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 850930), an NSF postdoctoral fellowship, and an NSERC Discovery Grant. The 4th author is grateful to the Institute for Advanced Study, and Bob Moses for funding her membership there.
@@ 2. Background
Assume throughout this section that $\Gamma$ is a finite oriented graph where each vertex has valence at least $3$ and $F_{r}$ is a free group of rank $r\geq 3$.
@@ 2.1. Edge Maps on Graphs
Suppose $\Gamma$ has positively oriented edges $\{e_{1},e_{2}\dots,e_{n}\}$ and vertices $\{v_{1},v_{2},\dots,v_{m}\}$. We use the notation $E\Gamma:=\{e_{1},e_{2}\dots,e_{n}\}$, and $V\Gamma:=\{v_{1},v_{2},\dots,v_{m}\}$, and $E^{\pm}\Gamma=\{e_{1},\overline{e_{1}},\dots,e_{n},\overline{e_{n}}\}$, with an overline indicating a reversal of orientation. Given $v\in V\Gamma$, a direction at $v$ will mean an element of $E^{\pm}\Gamma$ with initial vertex $v$. We let $\mathcal{D}\Gamma$ denote the set of directions at vertices in $\Gamma$. A turn at $v$ will mean an unordered pair $\{d_{1},d_{2}\}$ of directions at $v$. The turn is degenerate if $d_{1}=d_{2}$.
An edge path (or path) $\rho$ in $\Gamma$ is a finite sequence $(a_{1},a_{2},\dots,a_{\ell})\in(E^{\pm}\Gamma)^{\ell}$ such that there exists a sequence $(v_{1},v_{2},\dots,v_{\ell-1})\in(V\Gamma)^{\ell-1}$ satisfying that the turn $\{\overline{a_{j}},a_{j+1}\}$
[Uncaptioned image]is a turn at $v_{j}$ for each $j\in\{1,2,\dots,\ell-1\}$. For such a path $(a_{1},a_{2},\dots,a_{m})$ we write $\gamma=a_{1}a_{2}\dots a_{m}$ and say $\gamma$ contains the oriented edges $a_{1},a_{2},\dots,a_{m}$ and takes the turns $\{\overline{a_{1}},a_{2}\}$, $\{\overline{a_{2}},a_{3}\}$, $\dots$, $\{\overline{a_{n-1}},a_{n}\}$. We call $\gamma$ tight if it takes no degenerate turns, which we colloquially describe as there being no “backtracking.”
@@
To $g$ we associate a direction map $Dg:\mathcal{D}\Gamma\to\mathcal{D}\Gamma^{\prime}$ such that if $g(e)=a_{1}a_{2}\dots a_{m}$, for some $m\geq 1$ and $a_{1},a_{2},\dots,a_{m}\in E^{\pm}\Gamma$, then $Dg(e)=a_{1}$. Now suppose that $g\colon\Gamma\to\Gamma$ is a self-map. We call a direction $e$ periodic if $Dg^{k}(e)=e$ for some $k>0$, and fixed if $k=1$. The turn $\{d_{1},d_{2}\}$ is called an illegal turn for $g$ if $\{Dg^{k}(d_{1}),Dg^{k}(d_{2})\}$ is degenerate for some $k$. Defining an equivalence relation on $\mathcal{D}\Gamma$ by $d_{1}\sim d_{2}$ when $\{d_{1},d_{2}\}$ is an illegal turn, the equivalence classes are called gates. Note that each gate at each periodic vertex contains a unique periodic direction.
Viewing $g$ as a continuous map of graphs, $g$ represents $\varphi$ when $\pi_{1}(\Gamma)$ has been identified with $F_{r}$ (that is $\Gamma$ is marked) and $\varphi$ is the induced map of fundamental groups. When a marking is not explicitly given, we mean “there exists a marking such that.”
@@ 2.2. Train track maps and fully irreducible outer automorphisms
A nonnegative integral matrix $A=[a_{ij}]$ is irreducible if for each $(i,j)$, there is a $k\in\mathbb{Z}_{>0}$ so that the ${ij}^{th}$ entry of $A^{k}$ is positive, and so in particular is at least $1$. If $A^{k}$ is strictly positive for some $k\in\mathbb{Z}_{>0}$ then $A$ is primitive. Furthermore, $A$ is Perron–Frobenius (PF) if there exists an $N$ such that, for each $k\geq N$, we have that $A^{k}$ is strictly positive.
For nonnegative integral matrices, being primitive is equivalent to being irreducible plus aperiodic. If $M=M(g)$ is primitive, then by Perron–Frobenius theory, $M$ has a unique eigenvalue $\lambda(g)$ of maximal modulus. $\lambda(g)\in\mathbb{R}_{>1}$ and is the Perron–Frobenius (PF) eigenvalue of $M$. It is called the stretch factor (or dilatation) of $g$.
A tt map is irreducible if its transition matrix is irreducible. Not every element of $\textup{Out}(F_{r})$ is represented by a tt map, and even fewer by irreducible tt maps. An outer automorphism $\varphi\in\textup{Out}(F_{r})$ is fully irreducible if no positive power preserves the conjugacy class of a proper free factor of $F_{r}$. Bestvina and Handel [BH92] proved that each fully irreducible outer automorphism admits expanding irreducible tt representatives and that the stretch factor of these representatives is an invariant of the outer automorphism.
By [FH11, Corollary 4.43], for each $r\geq 2$, there exists an $R(r)\in\mathbb{N}$ such that for each expanding irreducible tt representative $g$ of a fully irreducible $\varphi\in\textup{Out}(F_{r})$, among other properties, each periodic direction is fixed by $g^{R(r)}$. This power $R$ is called the rotationless power.
@@ 2.3. Whitehead graphs & lamination train track (ltt) structures
Let $g:\Gamma\to\Gamma$ be a tt map. The local Whitehead graph $\operatorname{LW}(g;v)$ at a $v\in V\Gamma$ has a vertex for each direction at $v$ and an edge connecting the vertices corresponding to a pair of directions $\{d_{1},d_{2}\}$ at $v$ precisely when the turn $\{d_{1},d_{2}\}$ is $g^{k}$-taken for some $k\in\mathbb{Z}_{>0}$. Given a fixed vertex $v$, the stable Whitehead graph $\operatorname{SW}(g;v)$ is the restriction of $\operatorname{LW}(g;v)$ to the periodic direction vertices and the edges connecting them. In terms of gates, $\operatorname{SW}(g;v)$ has a vertex for each gate at $v$.
@@
$\operatorname{IW}(\varphi)=\bigsqcup_{v\in V\Gamma}\operatorname{SW}(g;v),$
@@
The lamination train track (ltt) structure $G(g)$ is obtained from its underlying graph $\Gamma$ by replacing each vertex $v\in V\Gamma$ with $\operatorname{LW}(g;v)$: replace $v$ with a vertex for each directed edge at $v$ labeled with that direction. Then identify each of these new vertices with the corresponding vertex of $\operatorname{LW}(g;v)$. Vertices and edges of $\operatorname{SW}(g;v)$ are colored purple and the remaining vertices and (open) edges are colored red. Alternatively, one could start with $\bigsqcup_{v\in V\Gamma}\operatorname{LW}(g;v)$, color $\operatorname{LW}(g;v)$ as just described, and then include a directed edge $[e,\overline{e}]$ for each directed edge $e\in E\Gamma$. To simplify figures, if the local Whitehead graph at a vertex of $\Gamma$ is complete we do not always draw it.
Figure 1. The right-hand image is the ltt structure $G(\mathfrak{g})$ for the map $\mathfrak{g}$ of (11) with the taken turns as in (12). The 3 local Whitehead graphs are colored, with the stable Whitehead graphs colored purple (missing only $\bar{c}$). The ideal Whitehead graph is the union of the purple graphs, i.e. of the stable Whitehead graphs.